OpenStudy (anonymous):

solve x^4-34x^2=-225

OpenStudy (anonymous):

You can do a variable change: \[y = x^2\]

OpenStudy (anonymous):

then, \[y^2 - 34y + 225 = 0\]

OpenStudy (anonymous):

Can you do the rest by yourself?

OpenStudy (anonymous):

quadratic equation?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

so a=1 b=34 c=225

OpenStudy (anonymous):

exactly!. Once you get the two values of y,you return to y = x^2 and solve for x

OpenStudy (anonymous):

k thank you

OpenStudy (anonymous):

i got a giant number :(( can you draw it out for me

OpenStudy (anonymous):

well i got \[y_1 = 25, y_2 = 9\]

OpenStudy (anonymous):

But, y = x^2, therefore, \[x_1 = 5, x_2 = -5, x_3 = 3, x_4 = -3\]

OpenStudy (anonymous):

do you think you can draw it out?

OpenStudy (anonymous):

Draw it?

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):



OpenStudy (anonymous):

That was rude, i understand draw. :-S